Need help with a rate of change of function problem. G(x)=x^3-2x+1
From -3 to -2
From -1 to 1
From 1 to 3 I'm just not sure how to set these problems up, and how to work them! Hope you can help! Thank you!

Answer by Alan3354(69443)

Need help with a rate of change of function problem. G(x)=x^3-2x+1
From -3 to -2
From -1 to 1
From 1 to 3
-------------
G(x)=x^3-2x+1
From -3 to -2
---------
Find G(-3) & G(-2)
G(-3) = (-3)^3 - 2*(-3) + 1 = -27 + 6 + 1 = -20
G(-2) = (-2)^3 - 2*(-2) + 1 = -8 + 4 + 1 = -3
-------
ROC = (-3 - (-20))/(-2 - (-3))
= 17/1
= 17
===========
Do the others the same way.